OpenStudy (anonymous):

what are the solutions to 3x^2=8x-6?

OpenStudy (anonymous):

3x^2=8x-6 3x^2-8x+6=0 A B C -B+/_ the sq root of B^2-4(A)(C) divided by 2(A) Now we have +8+/_ sq root 8^2-4*3*(-6) divided by 2*3 8+/_sq rt 136/6 = 8+/_4sqrt34/3 reduced to 4+/_2sq rt34/3 and as two answers one will be \[4+2\sqrt{34}/3 and 4-2\sqrt{34}/3\]